Circular No. 6118

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1988ad IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of an

apparent supernova (V = 19.5) on two technical pan plates obtained by J. Heudier and himself with the OCA Schmidt telescope on 1988 Dec. 28.77 and 31.79 UT. SN 1988ad is located at α = 1^{h}22^{m}50^{s}.65, δ = −1^{o}38′26′′.1 (equinox 1950.0), which is 5′′.7 west and 0′′.6 north of the center of an S0 galaxy numbered 36 by Dressler in his study of the cluster Abell 194 (1980, Ap.J. Suppl. 42, 565). A nearby star (mag ∼ 16.5) is at position end figures 47^{s}.58, 37′07′′.6. SN 1988ad does not appear on the digitized SRC J survey, although the background brightness of the galaxy is high in this vicinity. Nothing is present at the position of SN 1988ad on films of similar resolution and limiting magnitude (21–22) obtained in 1994 December. Pollas notes that the host galaxy is not a member of Abell 194 since its heliocentric radial velocity is 20 269 km/s (Lucey and Carter 1988, MNRAS 235, 1177).

GRO J0422+32 C. Chevalier and S. A. Ilovaisky, Haute-Provence Observatory, commu-

nicate: “CCD photometry of this x-ray transient’s quiescent optical counterpart has been obtained with the 1.2-m Haute-Provence telescope during 8 nights between Oct. 30 and Dec. 6, mainly in the Cousins R band. The periodogram of these data shows two peaks: one at 0.106133 day, and the other at twice this value, 0.212265 ± 0.000072 day (5.0944 ± 0.0017 hr). This latter value excludes that reported by Orosz and Bailyln (IAUC 6103). This new period could fit the Haute-Provence data obtained during the active state, and it differs by only 0.06 percent our previous result (A.Ap., in press). The 5.094-hr R-band lightcurve is double-waved (maximum peakto-peak amplitude 0.2 mag) and shows variations in shape and amplitude superposed onto night-to-night variations in mean brightness. The average magnitudes for the counterpart are R_{c} = 21.03 ± 0.10, V = 21.96 ± 0.20, I_{c} = 19.88 ± 0.15.”

PERIODIC COMET BORRELLY (1994l) Total visual magnitude estimates: Nov. 28.25 UT, 7.8 (C. E. Spratt,

Victoria, BC, 0.20-m reflector); 30.93, 8.4 (K. Sárneczky, Ráktanya, Hungary, 20×60 binoculars); Dec. 4.44, 7.9 (J. E. Bortle, Stormville, NY, 10×50 binoculars); 11.12, 7.9 (S. Foglia, Lumezzane, Italy, 0.40-m reflector).

1994 December 23 Daniel W. E. Green